A tank can be filled by one pipe stream in 6 hours, and by another pipe stream in 4 hours. It can be emptied by third pipe stream in 7 hours. If the first pipe stream was opened two hours ahead of the second and third pipe streams, how long will it take to fill the tank?
Answer by mananth(16949)

first pipe = 6 hours
so it can do 1/6 of the job in 1 hour.
so in 2 hours it has done 1/3 of the of the job.
Balance job remaining = 1-1/3 = 2/3 of the job.
...
Time taken by three pipes to complete balance job
pipe 1- 4 hours
pipe 2- 8/3 hours
pipe 3 = 14/3 hours
...
1/4+3/8 -3/14 of the job they will do in 1 hour
0.41 of the job they do in1 hour.
1/0.41 hours they will take to do 2/3 of the job
=2.4 hours
+ 2hours (pipe one alone)
= 4.4 hours
...
m.ananth@hotmail.ca